According to a Pew Research Center study, in May 2011, 44% of all American adults had a smart phone (one which the user can use to read email and surf the Internet). A communications professor at a university believes this percentage is higher among community college students. She selects 437 community college students at random and finds that 195 of them have a smart phone. Then in testing the hypotheses:

H0: p = 0.44 versus

Ha: p > 0.44,

what is the test statistic?

z =  . (Please round your answer to two decimal places.)
Answer #1

\hat p = x / n = 195/437=0.4462

P0 = 0.44

1 - P0 = 1 - 0.44=0.56

Test statistic = z

= \hat p - P0 / [\sqrtP0 * (1 - P0 ) / n]

=0.4462 - 0.56 / [\sqrt(0.44*0.56) / 437]

z=0.262

z = 0.26